Spreadsheet 1: Amortization Table

Create an amortization table in MS-Excel in the format shown below:

LOAN ANALYSIS: Prepared by Kent Saunders Principal $10,000.00 Annual interest 10.00% 10 Term (years) Payment periods per year No. of payments Periodic rate of interest 5.00% Basic P ayment per period$802.43 Added payment Pmt Beg. Principal Added AppliedEnding Cumulativie Balance Interest Payment Payment Principa Balance Interest 500.00 9,697.57 484.88802.430.00 317.559,380.02984.8S 9,380.02469.00 802.43 0.00 333.43 9,046.591,453.88 9.04659 52.33 802.43 0.00 350.10 8.696.49 906.21 10,000.00 500.00802.430.00 302.43 9,697.57

Scenario: 2 years ago Janice got a $100,000, 15-year mortgage with an annual interest rate of 6% and monthly payments.

1) What is her monthly payment?

2) How much does she owe today (after 24 payments)?

3) How much will she owe in 3 years (after 60 payments)?

4) How much will she owe in 3 years (after 60 payments) if she makes an extra $200 payment every month starting today?
Answer #1
1) Monthly payment using loan amortization formula = 100000*0.005*1.005^180/(1.005^180-1)= $             843.86
2) Amount owed after 24 payments = PV of the remaining 156
installments = 843.86*(1.005^156-1)/(0.005*1.005^156) = $       91,255.39
3) Amount owed after 60 payments = PV of the remaining 120
installments = 843.86*(1.005^120-1)/(0.005*1.005^120) = $       76,009.38
4) CV of 200 paid for 60 months = 200*(1.005^60-1)/(0.005) = $       13,954.01
Amount owed at the end of 60 months = 76009.38-13954.01 = $       62,055.38